OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? A. (2, −3) B. (−2, 1) C. (5, −2) D. (−6, 5)

OpenStudy (anonymous):

it is A

OpenStudy (anonymous):

plug in 2 for x and -3 for y

OpenStudy (anonymous):

ok can you explain why it is A?

OpenStudy (anonymous):

We can re-order the first equation: x + y = -1 Subtract 1 to both sides: x = -y - 1 Now we can plug in -y - 1 for 'x' in the 2nd equation. x - 3y = 11 -y - 1 - 3y = 11 Simplify: -4y - 1 = 11 Can you finish that?

OpenStudy (anonymous):

@Glorm321 Don't give answers. Read the Code of Conduct: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

no sorry not really good at math anymore

OpenStudy (anonymous):

so two plus negative three is negative 1

OpenStudy (anonymous):

2 minus negative three times negative three equals 2+9 which equals 11

OpenStudy (anonymous):

thanks that really helps : )

OpenStudy (anonymous):

and I am actually new so I did not know about the code of conduct! @iGreen.

OpenStudy (anonymous):

do you understand the problem?

OpenStudy (anonymous):

its cool at least you explained it for me and yes thank you can you help with a few more?

OpenStudy (anonymous):

@Glorm321 Yes, I can see that. But please, in the future, guide them to the answer so that they get it on their own. Don't give them the answer AT ALL.

OpenStudy (anonymous):

lol and i'm new too i just started yesterday

OpenStudy (anonymous):

well I will give you a tip @TeenWolfGirl , when you are given a problem with coordinates, plug them into the equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i don't mean to be a bother but can you, one more time, explain it plz

OpenStudy (anonymous):

@iGreen can you plz explain it again?

OpenStudy (anonymous):

Okay, so we first rearrange the first equation so that we can substitute. x + y = -1 For us to substitute we have to get one of the variables on one side by itself. So we can subtract 'y' to both sides. x = -y - 1 So now we can plug in -y - 1 into the 2nd equation: x - 3y = 11 -y - 1 - 3y = 11

OpenStudy (anonymous):

Then we simplify it and solve for 'y'.

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

kinda sorry i am taking so like i am writing this down so that i an study it

OpenStudy (anonymous):

Okay, no problem.

OpenStudy (anonymous):

Our next step would be to simplify '-y - 3y', which is the same thing as '-1 - 3', then we add a 'y' after it. So that gives us -4y. Plug that back in: -4y - 1 = 11 Now we add 1 to both sides to get the variable by itself. -4y = 12 Then, our last step is to divide -4 to both sides, which gives us: y = -3

OpenStudy (anonymous):

Now we have the y-value of the solution. We have to plug in y = -3 into any of the two original equations to find the x-value. x + y = -1 Plug in -3 for y: x - 3 = -1 Add 3 to both sides: x = 2 So x = 2, so our final answer is x = 2, y = -3, or (2, -3).